自然言語理解 2 単位 (選択)

Natural Language Understanding  青江順一 · 教授/ システム創生工学専攻知能情報システム工学コース知能工学講座

泓田正雄 · 講師/ システム創生工学専攻知能情報システム工学コース知能工学講座，森田和宏 · 講師/ システム創生工学専攻知能情報システム工学コース基礎情報工学講座

【授業目的】本講義の目的はヒューマンコミュニケーションと知的テキスト検索 ; アプリケーションに対する自然言語理解手法を講述する．特に，構文解析ア ; ルゴリズムと意味解析技術を種々のアプリケーションを対象として説明する． ; 【授業概要】自然言語理解はヒューマンコミュニケーションと知的テキスト検索 ; 分野の最も重要な技術である．この技術は知識辞書と解析と結びつける効率 ; 的な連携に依存する．この講義では，知識を使用する意味情報と格解析を融 ; 合した格構造について説明する．そして，最先端の自然言語アプリケーショ ; ンシステムについても具体例で説明し，演習課題を導入する．演習課題は，知 ; 識辞書のトライ構造，形式的文書記述と理解，特許文書と理解などを含む．本 ; 科目は，工業に関する科目である． ; 【授業形式】講義形式とポートフォーリオ形式の併用 ; 【キーワード】対話理解，意図理解，知識辞書，形態素解析，特許文書 ; 【先行科目】『自然言語処理』 (1.0)，『データマイニング』 (1.0) ; 【関連科目】『機械翻訳特論』 (0.5)，『言語モデル論』 (0.5) ; 【履修要件】 [要件] ; 【履修上の注意】 [注意] ; 【到達目標】 ; 1. 自然言語理解，対話理解，意図理解の技術を習得できる． ; 2. 自然言語処理におけるキー検索技法を習得できる． ; 3. 形態素解析，仮名漢字変換の技術を習得できる． ; 4. 言語理解に基づく技術文書，特許文書の記述方法を習得できる． ; 【授業計画】 ; 1. 自然言語理解の概要 ; 2. 対話理解 ; 3. 意図理解 · レポート ; 4. 作文における新情報 ; 5. 作文構造の設計 · レポート ; 6. 知識辞書 ; 7. キー検索アルゴリズム ; 8. トライ構造 · レポート ; 9. 形態素解析 ; 10. 仮名漢字変換 · レポート 11. 技術文書の理解

12. 技術文書の理解演習 · レポート

13. 技術文書の理解

14. 特許文書の理解

15. 特許文書理解に関する演習 · レポート

16. レポート

【成績評価基準】出席 (20%)，レポート (80%) として評価する．

【教科書】資料を配布する【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216684 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 青江 (Dr. 棟 604, 088-656-7486, aoe@is.tokushima-u.ac.jp) ���� (オフィスアワー: 木曜日 14:00∼ 17:00(年度ごとに学科の掲示を参照すること))

Natural Language Understanding 2 units (selection)

Jun-ichi Aoe · Professor / Intelligent Systems, Information Science and Intelligent Systems, Systems Innovation Engineering , Masao Fuketa · Associate Professor / Intelligent Systems, Information Science and Intelligent Systems, Systems Innovation Engineering Kazuhiro Morita · Associate Professor / Information Science, Information Science and Intelligent Systems, Systems Innovation Engineering

Target⟩ This lecture is focused on methods of Natural Language Understanding for human-communication and intelligent text search applications. Particulary, syntax analysis algorithms and semantic analysis technics for various kinds of applications are explained in this lecture.

Outline⟩ Understanding is most important technics for the human-communication and the intelligent text search. Efficiency of the technics depend on the relationsips between knowledge dictionaries and analysis algorithms. In this lecture, case-structure with semantic informations and case-analysis algorithms is explained. Moreover, the newest Natural Language Processing systems are introduced with detailed examples. Exercises include Trie structure for the semantic dictionary, understanding for formal document, and understanding for patent documents. This subject of lecture belongs to industrial field.

Style⟩ Lecture in combination with Portfolio �������⟩ dialogue understanding, intention understanding, knowledge dictio-

nary, morphological analysis, patent document Fundamental Lecture⟩ “自然言語処理”(1.0), “データマイニング”(1.0) Relational Lecture⟩ “Advanced Machine Translation”(0.5), “Language Modeling”

(0.5)

Requirement⟩ [要件] Notice⟩ [注意]

Goal⟩

1. Technics of Natural Language Understanding, Dialogue Undierstanding and Intention Understanding can be learned.

2. Key search methods in Natural Language Processing can be learned.

3. Technics of Morphological Analysis and Kana-Kanji Translation can be learned.

4. How to describe technical document and patent document based on Language

Understanding can be learned.

Schedule⟩

1. Introduction

2. Dialogue Understanding

3. Intention Understanding, Report 4. New Information on Composition

5. Design of Composition Structure, Report

6. Knowledge Dectionary

7. Key Search Algorithm

8. Trie Structure, Report

9. Morphological Analysis

10. Kana-Kanji Translation, Report

11. Understanding of Technical Document

12. Exercise for Understanding of Technical Document, Report

13. Understanding of Technical Document

14. Understanding of Patent Ducument

15. Exercise for Understanding of Patent Document, Report

16. Report

Evaluation Criteria⟩ Result is decided by attendance(20%) and reports(80%).

Textbook⟩ Printed synopsises of lecture are distributed.

Reference⟩ [参考資料]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216684

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ Aoe (Dr604, +81-88-656-7486, aoe@is.tokushima-u.ac.jp) ���� (Office Hour: 木曜日 14:00∼ 17:00(年度ごとに学科の掲示を参照すること))